# 4DHands: Reconstructing Interactive Hands in 4D with Transformers

### v2

@@ 4.1 Implementation details
We implement our model based on PyTorch [38]. We use a pretrained Vision Transformer following a huge design (ViT-h) as the backbone to encode image features. The resolution of input image is $256\times 192$ and the size of the feature map is $H=16,W=12,C=1280$. We use transformer encoders and decoders in our feature fusion modules. We train our model using AdamW optimizer with a linear learning rate decay schedule from an initial learning rate $1\times 10^{-5}$. The length of an image sequence is set to 9 with a gap of 5 frames to sample images in a consecutive video sequence.
@@ 4.2 Datasets
Interhand2.6M [34]is the first publicly available dataset of sequences of interacting hands with accurate annotations, which is widely used for training and testing in two-hand reconstruction methods. InterHand2.6M in 30 fps is for video-related research and the 5 fps version is a subset of the 30 fps version which removes the redundancy for single image research. In our method, we make use of 30 fps and 5 fps interacting two-hand (IH) data with human and machine (H+M) annotations and filter invalid data following [27]. We ultimately utilize 1,614K images for training and 601K images for testing.
@@ 4.3 Metrics
We adopt Mean Per Joint Position Error (MPJPE) and Mean Per Vertex Position Error (MPVPE) and Area Under Curve (AUC) scores to measure the accuracy of the pose and shape of the estimated meshes, and Mean Relative-Root Position Error (MRRPE) to measure the accuracy of relative positions of two hands. Following the prior method [27], we perform root joint alignment and skeleton scaling when evaluating MPJPE. In addition, we use acceleration error Accel_E to evaluate the temporal consistency, following [22].
@@ 4.4 Qualitative Results
Our qualitative comparisons are shown in Figure 3. We demonstrate results of in-the-wild datasets ARCTIC and RenderIH, which are unseen in the training phrase. We compare with IntagHand[27] and ACR [53] in camera perspective and an arbitrary perspective to show more details. We show more results on realistic complex cross-hand interacting samples in 4. The results demonstrate the stability of 4D-Hands in real-world scenarios.
Figure 3: Qualitative Comparison. We compare 4DHands with state-of-the-art methods on in-the-wild datasets ARCTIC[10] and RenderIH[26]. ARCTIC provides two-hand object interacting images, and RenderIH provides interactive hands under different lighting conditions. We have demonstrated the results from both the camera perspective and an arbitrary perspective. The results show that our method outperforms the state-of-the-art methods in complex real-world environments.
Figure 4: Qualitative Results. We show our model’s results on complex cross-hand interactions in realistic scenarios to demonstrate its performance. For a detailed perception, views of multiple perspectives are demonstrated.
@@ 4.5 Comparison with state-of-the-art methods
To demonstrate the versatility and stability of 4DHands across various scenarios, we conducted comparative analyses with state-of-the-art methods in multiple domains, including temporal two-hand evaluation, frame-based two-hand evaluation, temporal single hand evaluation and in-the-wild dataset evaluation.
Comparison with Temporal Two Hand Methods. In Table 1 we compare the proposed 4D-Hands with previous temporal methods on Interhand2.6m 30fps. The results show that our method far outperforms the existing two-handed temporal models, with 45.20%(7.37 mm $vs.$ 13.45 mm) improvement in MPJPE and 24.05% (2.81 mm/$s^{2}$ $vs.$ 3.70 mm/$s^{2}$) improvement in Accel_E.
Comparison with Frame-based Two Hand Methods. To compare with state-of-the-art frame-based methods and demonstrate the performance in single frame input scenarios, we conduct experiments at Interhand2.6m 5fps by expanding single image to sequences. As can be observed at Table 2, even though we drop the temporal knowledge, our method still keep high performance and outperforms previous methods that were specially designed for single frame input scenarios, achieving 7.49 mm, 7.72 mm and 24.58 mm in MPJPE, MPVPE and MRRPE.
Table 1: Quantitative comparisons with frame-based and temporal methods on Interhand2.6m 30fps dataset.
@@
Table 2: Performance comparisons on the Interhand2.6m 5fps dataset. We compare with the most recent methods.
@@
Comparison with Temporal Single Hand Methods. To further exploit the universality of our method, we conduct experiments on sequential single hand dataset Dex-YCB. We manage our input by flipping an input image as the opposite hand, and using far away bounding boxes to indicate the hands are not interacting. We report the results of Root-Aligned and Procrustes-aligned metrics. It can be observed in Table 3 that our method outperforms all existing single hand temporal methods in all metrics. It is noteworthy that our lead in Root-Aligned metrics is greater than those in Procrustes-aligned metrics (+38.63% $vs.$ +9.00% in MPJPE, +12.70% $vs.$ +0.89% in JAUC), indicating that we can predict accurate shapes and rotation angles without the assistance of ground truth alignment.
@@
Comparison on In-the-wild Dataset. We report the results on HIC dataset, which contains in-the-wild cross-hand data, to compare the performance at in-the-wild scenarios. We only use samples in which both hands are visible. In 4 we compare the results with IntagHand and InterWild, a SOTA method specifically designed for two-hand recovery in the wild. Comparing to the MPJPE results on Interhand2.6m, which is employed during training, the performance of IntagHand and InterWild deteriorates by 131.8% and 40.5%, respectively, while our method exhibits a more modest reduction of 24.43%. This demonstrates that our method has greater stability on unseen data.
Table 4: Performance comparisons on the HIC dataset. We compare with Intaghand [27] and InterWild [33].
[…]
Intaghand[27] 20.38 21.56 73.04
InterWild [33] 15.62 15.17 26.43
Ours 9.32 9.93 25.26
@@ 4.6 Ablation Study
ViT Baseline. We trained a baseline using ViT and a transformer decoder on full training set, cross-hand feature interaction is not employed on the baseline. We show the results in 5. Although ViT has high performance in detecting single hand, the occlusion caused by hands interaction reduces its predictive ability. Besides, without cross-hand knowledge, it separately predict positions of two hand, causing huge errors in relative position.
@@
Effectiveness of Cross-hand Information. We experimented in adding cross-hand information in two ways: (1) Input holistic image including two hands(+Cross H). (2) Input separate images, each image contains a single hand(+Cross S). In both way we add attention transformer module to integrate cross-hand information, the results show that the last way achieves better accuracy. The main reason is that by separating two hands as input, the visual encoder focuses on extracting single hand features, allowing the ViT to learn more hand details when learning on large scale datasets.
@@
Effectiveness of Relation-aware Tokenization. By removing RAT(+Cross S w/o RAT), the model lost global position information during inference, which means that the model could only infer two-hand relative position from visual information. In InterHand2.6m, the two hands are relatively close and the 2D scales of each hand are similar. Therefore, even without a global position information, the approximate positions of both hands can be inferred from the input image. We test cases when two hands are relatively far away and cases when 2D scales of each hand are different. As shown in 5, without RAT, the model performs poor in predicting two-hand relative position.
@@
Effectiveness of Temporal Fusion. By adding temporal information interacting(+Cross S + Seq), our method can make sequential consistent prediction therefore infer reasonable poses encountering occlusions, and achieves improvement of 0.32 mm in MPJPE. In Figure 6 we visualize the predicted mesh sequences in the case of severe occlusion. It shows that the left hand is entirely occluded, the model without temporal fusion makes erroneous prediction. Our full model predicts a consistent sequence with assistant information from the past and the future.
Table 5: Ablation studies of the RAT and SIR modules on InterHand2.6M.

### v3

@@ 3.4.3 Inter-Hand Losses
3.4.4 Sequential Consistency Losses
When using video as input, we use a smooth loss function to maintain the smoothness of the predicted mesh between video frames and reduce jitter:
$L_{consis}=\sum_{t=1}^{N_{t}-1}||\theta_{t+1}-\theta_{t}||^{2}$ (16)
where $N_{t}$ is the length of video sequence.
When taking multi-view images as input, the inter-hand relative distance and relative rotation angle should be identical in all input images. Since the extrinsic parameters of the cameras from each viewpoint are different, predictions of the rotation angles and relative distances for each are coupled with the camera’s rotation and translation predictions. Therefore, in each frame, we perform the same rotation and translation operations on the predicted joints of both hand to eliminate the effect of camera parameters. This process moves the root joint of the right hand to the coordinate origin and sets the rotation angle to zero, then denote the transformed left hand joints as $J^{*}$. The consistency loss is calculated:
$L_{consis}=\sum_{i=1}^{N_{t}}\sum_{k=1}^{N_{t}}||J^{*}_{i}-J^{*}_{k}||^{2}$ (17)
@@ 4.1 Implementation details
We implement our model based on PyTorch [43]. We use a pretrained Vision Transformer following a huge design (ViT-h) with 32 layers as the backbone to encode image features. The resolution of input image is $256\times 192$ and the size of the feature map is $H=16,W=12,C=1280$. We use transformer encoders and decoders in our feature fusion modules. We train our model for 1,000,000 steps using AdamW optimizer with a linear learning rate decay schedule from an initial learning rate $1\times 10^{-5}$, and the decay rate is 0.7 for every 200,000. The training set is mixed of Interhand2.6m [39], Re:Interhand [41] and Single Hand Datasets, we use random sampling strategy with sampling rate 40%, 40%, 20%. When taking temporal sequence as input, the length of an image sequence is set to 9 with a gap of 5 frames to sample images in a consecutive video sequence.
@@ 4.2 Datasets
Interhand2.6M [39]is the first publicly available dataset of sequences of interacting hands with accurate annotations, which is widely used for training and testing in two-hand reconstruction methods. InterHand2.6M in 30 fps is for video-related research and the 5 fps version is a subset of the 30 fps version which removes the redundancy for single image research. In our method, we make use of 30 fps and 5 fps interacting two-hand (IH) data with human and machine (H+M) annotations and filter invalid data following [32]. We ultimately utilize 1.6M images for training and 1.2M images for testing.
@@ 4.3 Metrics
We adopt Mean Per Joint Position Error (MPJPE) and Mean Per Vertex Position Error (MPVPE) and Area Under Curve (AUC) scores to measure the accuracy of the pose and shape of the estimated meshes, and Mean Relative-Root Position Error (MRRPE) to measure the accuracy of relative positions of two hands. For mean position error and AUC we get results under both Root-relative analysis (Ra) and Procrustes analysis (Pa) .Following the prior method [32], we perform root joint alignment and skeleton scaling when evaluating MPJPE. In addition, we use acceleration error Accel_E to evaluate the temporal consistency, following VIBE[27].
@@ 4.4 Qualitative Results
For monocular inputs, our qualitative comparisons are shown in Figure 3. We demonstrate results of in-the-wild datasets ARCTIC [13] and RenderIH [31], which are unseen in the training phrase. We compare with IntagHand[32] and ACR [64] in camera perspective and an arbitrary perspective to show more details. We show more results on realistic complex cross-hand interacting samples in Figure 4. The results demonstrate the stability of OmniHands in real-world scenarios.
For multi-view inputs, the qualitative results are shown in Figure 5, where 4 camera views are selected as demonstration. In every perspective, the hands are obscured to varying degrees. It can be observed that by integrating multi-view information, the hand interaction information is mutually supplemented, allowing successful reconstruction of the hand mesh corresponding to the respective view, even when the hands are completely occluded in some perspectives. It is worth noting that due to the absence of camera parameter inputs, our model needs to predict the camera angle solely based on visual information and determine the hand orientation in the presence of occlusion, which is a non-trivial achievement.
Figure 3: Qualitative Comparison. We compare OmniHands with state-of-the-art methods on in-the-wild datasets ARCTIC[13] and RenderIH[31]. ARCTIC provides two-hand object interacting images, and RenderIH provides interactive hands under different lighting conditions. We have demonstrated the results from both the camera perspective and an arbitrary perspective. The results show that our method outperforms the state-of-the-art methods in complex real-world environments.
Figure 4: Qualitative Results of in-the-wild hard cases. We show our model’s results on complex cross-hand interactions in realistic scenarios to demonstrate its performance. For a detailed perception, views of multiple perspectives are demonstrated.
@@ 4.5 Comparison with state-of-the-art methods
To demonstrate the versatility and stability of OmniHands across various scenarios, we conducted comparative analyses with state-of-the-art methods in multiple domains, including temporal two-hand evaluation, frame-based two-hand evaluation, temporal single hand evaluation and in-the-wild dataset evaluation.
Comparison with Temporal Two Hand Methods. In Table 1 we compare the proposed OmniHands with previous temporal methods on Interhand2.6m 30fps [39]. To demonstrate the benefits of using large mixed training datasets for the model, as well as the impact of using temporal features on model accuracy, we trained four groups of models based on whether additional training data and temporal features were used. For the training datasets, we use Interhand2.6m [39], Re:Interhand [41], and Single Hand Datasets mixed up as our training set for a better model generalization. To ensure a fair comparison with other two-hand reconstruction methods, which are trained only on Interhand2.6m, we use the same training set to train our model and the results are also shown marked with †. As previously mentioned, OmniHands can take sequential frames or single frame as input, to figure out the impact of using temporal features, we report and compare the results of OmniHands with single frame and sequential frames as input, denoting single-frame version with §. The results show that our method far outperforms the existing two-handed temporal models, with 23.34%(7.37 mm $vs.$ 9.09 mm) improvement in MPJPE and 24.05% (2.81 mm/$s^{2}$ $vs.$ 3.70 mm/$s^{2}$) improvement in Accel_E. Comparing the results of whether using temporal features, it can be observed that using temporal feature results in better stability between frames and improvement of key points accuracy. The Accel_E improves 18.86% (2.81 mm/$s^{2}$ $vs.$ 3.34 mm/$s^{2}$) and 17.55% (3.02 mm/$s^{2}$ $vs.$ 3.55 mm/$s^{2}$) respectively, and there is also a notable improvement on MPJPE and MPVPE. By changing the training set, we can find that using extra training data can obviously improve the accuracy and stability.
Comparison with Frame-based Two Hand Methods. To compare with state-of-the-art frame-based methods and demonstrate the performance in single frame input scenarios, we conduct experiments at Interhand2.6m 5fps [39] by expanding single image to sequences. As can be observed at Table 2, even though we drop the temporal knowledge, our method still keep high performance and outperforms previous methods that were specially designed for single frame input scenarios, achieving 7.49 mm, 7.72 mm and 24.58 mm in MPJPE, MPVPE and MRRPE. We also report the results of multi-view frame inputs, without any camera parameter, our model achieves 5.93 mm, 6.20 mm and 14.08 mm in MPJPE, MPVPE and MRRPE, which significantly surpasses the accuracy of monocular estimation.
@@
Table 2: Performance comparisons on the Interhand2.6m 5fps dataset. We compare with the most recent methods. We also report the results of multi-view inputs, marked with $\diamond$.
@@
Ours $\diamond$ 5.93 6.20 14.08
Table 3: Multi-view single hand recovery results on Dex-YCB, comparing with other single hand multi-view methods.
[…]
MVP[69] 6.23 4.26 0.69 9.77 8.14 0.53
PE-Mesh-TR[63] 7.49 4.76 0.64 7.41 4.70 0.64
FTL-Mesh-TR[63] 8.66 5.51 0.59 8.75 5.75 0.59
POEM[63] 6.06 3.93 0.68 6.13 4.00 0.70
Multi-view Fit.[63] 7.33 5.29 0.65 7.22 5.19 0.65
Ours 6.44 3.18 0.87 6.42 3.26 0.87
[…]
Dataset Method Joints Vertices
[…]
Arctic[13] IntagHand[32] 24.88 12.02 0.57 0.76 26.30 12.52 0.55 0.75
InterWild[38] 30.21 14.75 0.50 0.71 31.49 15.14 0.48 0.69
DIR[47] 26.63 11.97 0.55 0.76 28.33 12.39 0.52 0.75
ours † 14.21 9.22 0.69 0.78 15.12 9.76 0.63 0.81
ours 13.57 7.98 0.74 0.84 14.51 8.75 0.72 0.82
HIC[55] IntagHand[32] 12.01 7.46 0.76 0.85 12.77 7.82 0.74 0.84
InterWild[38] 10.35 6.32 0.79 0.87 10.68 6.63 0.78 0.86
DIR [47] 11.90 7.12 0.76 0.85 12.79 7.55 0.75 0.84
ours † 9.47 5.55 0.81 0.88 10.35 6.07 0.79 0.87
ours 8.66 4.81 0.82 0.90 9.31 5.30 0.81 0.89
RenderIH[31] IntagHand[32] 33.84 15.19 0.47 0.70 35.17 15.34 0.45 0.69
InterWild [38] 15.92 10.54 0.69 0.79 16.50 10.59 0.68 0.78
DIR[47] 29.48 12.93 0.53 0.74 30.93 12.89 0.51 0.74
ours † 17.22 10.91 0.66 0.78 18.01 11.11 0.64 0.77
ours 13.71 9.17 0.72 0.81 14.32 9.48 0.71 0.81
Comparison with Temporal Single Hand Methods. To further exploit the versatility of our method, we conduct experiments on temporal single hand dataset Dex-YCB [4]. We manage our input by flipping an input image as the opposite hand, and using far away bounding boxes to indicate the hands are not interacting. We report the results of Root-Aligned and Procrustes-aligned metrics. It can be observed in Table 5 that our method outperforms all existing single hand temporal methods in all metrics. It is noteworthy that our lead in Root-Aligned metrics is greater than those in Procrustes-aligned metrics (+38.63% $vs.$ +9.00% in MPJPE, +12.70% $vs.$ +0.89% in JAUC), indicating that we can predict accurate shapes and rotation angles without the assistance of ground truth alignment.
@@
Quantitive Comparison on In-the-wild Datasets. We report the results on three in-the-wild datasets: Arctic [13], HIC [55] and RenderIH [31], to compare the performance at in-the-wild scenarios. Arctic contains images of two hand interacting with objects, HIC is a small dataset of two hand interaction sequences and RenderIH contains images of complicated two hand interactions under various light conditions. We only use samples in which both hands are visible. For fair comparison, temporal data or multi-view data are not used in these experiments, a single image is taken as input for each case. In Table 4 we compare the results with IntagHand [32], InterWild [38] and DIR [47]. IntagHand [32] and DIR [47] are two open-source inter-hand reconstruction models that have demonstrated excellent performance in training and testing on Interhand2.6m [39]. InterWild [38] is a SOTA method specifically designed for two-hand recovery in the wild. To demonstrate the efficiency of using large training data, we report results of models trained under two training set. In all three unseen datasets, our method achieves the best performance and leads in all metrics. It can also be found that by using large training set, our model can handle situations in more diverse scenarios, resulting in higher accuracy in all three testing sets. HIC [55] includes interactive gestures that are not present in Interhand2.6m. However, the scenes are relatively simple and the test samples are few, allowing other methods to also achieve good results. In HIC our method improves 16.33% (8.66 mm $vs.$ 10.35 mm) on Ra-MPJPE and 29.47% (4.81 mm $vs.$ 3.34 mm) on Pa-MPJPE. Arctic [13] includes a large number of images of two hands interacting with objects, which significantly differ from the data in Interhand2.6m. Additionally, there are many instances where parts of the hands are occluded, making the situation more complex. In this dataset, other methods performed poorly, whereas our method maintained robustness and achieved excellent results, improving 45.46% (13.57 mm $vs.$ 24.88 mm) on Ra-MPJPE and 33.61% (7.98 mm $vs.$ 12.02 mm) on Pa-MPJPE. In RenderIH [31], the diversity of lighting and backgrounds places high demands on the cross-scene robustness of the model. Our proposed Omnihands obtains 13.89% (13.71 mm $vs.$ 15.92 mm) improvement on Ra-MPJPE and 13.00% (9.17 mm $vs.$ 10.54 mm) on Pa-MPJPE.
Figure 5: Qualitive results of OmniHands with multi-view inputs on Interhand2.6m [39] and Arctic [13]. In each row, we present pairs of input and output images from 4 different camera views. The images on the left side of each pair serve as the multi-view sequence input.
Comparison on Single hand Multi-view Dataset. To demonstrate the performance of our model when taking multi-view images as input, in Table 3 we compare with other multi-view single hand methods on Dex-YCB [4]. We also compare with Multi-view pose estimation reported in [63]. It’s worth mentioning that other methods require camera parameters as input, which makes it easier for them to align different views but loses flexibility. Our method, on the other hand, does not need camera parameters as input. It only requires multiple images from different viewpoints, eliminating the tedious camera calibration steps while achieving higher accuracy than other methods.
@@ 4.6 Ablation Study
We evaluate the importance of each module in Omnihands by albation experiments in temporal input conditions, using temporal dataset Interhand2.6m 30fps as training set and tesing set. Moreover, we also conduct evaluations on the impact of camera numbers when taking multi-view images as input, using Interhand2.6m 5fps multi-view dataset.
4.6.1 Module Ablation
@@
ViT Baseline. We trained a baseline using ViT and a transformer decoder on full training set. Position map embedding, cross-hand feature interaction and temproal/multi-view interaction are not employed on the baseline, so the Bimanual Decoder, Collective Decoder and Individual Encoder are removed. We show the results in Table 6. Although ViT has high performance in detecting single hand, the occlusion caused by hands interaction reduces its predictive ability. Besides, without cross-hand knowledge, it separately predict positions of two hand, causing huge errors in relative position.
@@
Effectiveness of Cross-hand Information. We experimented in adding cross-hand information in two ways: (1) Input holistic image including two hands(+Cross H). (2) Input separate images, each image contains a single hand(+Cross S). In both way we add Bimanual Decoder module to integrate cross-hand information, the results show that the last way achieves better accuracy. The main reason is that by separating two hands as input, the visual encoder focuses on extracting single hand features, allowing the ViT to learn more hand details when learning on large scale datasets.
@@
Effectiveness of Relation-aware Tokenization. By removing position token map embedding in RAT(+Cross S w/o RAT), the model lost global position information during inference, which means that the model could only infer two-hand relative position from visual information. In InterHand2.6m [39], the two hands are relatively close and the 2D scales of each hand are similar. Therefore, even without a global position information, the approximate positions of both hands can be inferred from the input image. We test cases when two hands are relatively far away and cases when 2D scales of each hand are different. As shown in Table 6, without position token map, the model performs poor in predicting two-hand relative position.
@@
Effectiveness of Temporal Fusion. By adding Collective Decoder for temporal information interacting(+Cross S + Seq), our method can make sequential consistent prediction therefore infer reasonable poses encountering occlusions, and achieves improvement of 0.32 mm in MPJPE. In Figure 7 we visualize the predicted mesh sequences in the case of severe occlusion. It shows that the left hand is entirely occluded, the model without temporal fusion makes erroneous prediction. Our full model predicts a consistent sequence with assistant information from the past and the future.
4.6.2 Camera view Ablation
In multi-view reconstruction, the number of input camera views is important to the resulting accuracy. We tried to figure out the impact of camera views, and the experimental results are shown in Table 7. We divided the experimental results into two groups based on whether mixed training data was used. ’I’ indicates the use of only Interhand2.6m[39] as the training data, while ’I+R+A’ represents the use of a mixed training dataset comprising Interhand2.6m, Re:Interhand[41], and Arctic[13].
From the results presented in the table, it can be observed that simply switching from monocular input to 2-camera multi-view input significantly improves our prediction accuracy, and this improvement is even more pronounced when training with a mixed dataset. As the number of cameras increases from 2 to 16, the accuracy metrics improve accordingly regardless of the training set. This aligns with the intuitive understanding that more perspective leads to more accurate predictions. This demonstrates that our FIR module effectively utilizes 4D data to enhance prediction results and shows significant flexibility in handling 4D data of varying lengths. Additionally, by comparing the results of ’I’ and ’I+R+A’ in the table, it is evident that the accuracy with the mixed training dataset is significantly higher than with a single training dataset, irrespective of the number of cameras. This highlights our model’s capability to handle large-scale mixed data efficiently.
Table 6: Ablation studies of the RAT and FIR modules on InterHand2.6M.
@@ 4.6.2 Camera view Ablation
Table 7: Results on Interhand2.6m using different number of camera views. ”I” means only using Interhand2.6m training set, and ”I+R+A” means using Interhand2.6m, Re:Interhand and Arctic as training set.
[…]
I 2 7.02 7.28 20.59
4 6.88 7.13 18.90
8 6.79 7.05 17.91
12 6.65 6.93 16.86
16 6.49 6.76 17.08
I+R+A 2 6.51 6.74 17.72
4 6.28 6.55 16.49
8 6.21 6.47 16.26
12 6.14 6.42 14.15
16 5.93 6.20 14.08
5 Limitaions
Our method utilizes ViT-h as the backbone, which comprises 32 transformer layers, resulting in higher computational demands compared to other CNN-based methods. During inference, running on a single A800 GPU with Python code, our model achieves 16.05 FPS, indicating that it is not suitable for real-time rendering. Additionally, because we use external detection models to locate hands in images and regress the bounding boxes, the inference speed is also constrained by the performance of the detection models. We will focus on accelerating our model by reimplementing it in C++ instead of Python and optimizing the computation process to enable real-time rendering.